# A note on the Littlewood-Offord problem for discrete log-concave distributions

### v1

@@ 1 Introduction
Kleitman extended this result when $a_{1},\dots,a_{n}$ are vectors in a Hilbert space [14], [15]. Many variants of the Erdös-Littlewood-Offord problem have been established, such as improved bounds under certain constraints on the $a_{i}$’s [10], an inverse Littlewood-Offord theorem [26], [22], and a resilience version [2], but most are mainly dealing with Bernoulli distribution with parameter $1/2$. Recently, Fox, Kwan and Sauermann [8, Question 6.2] asked whether the Littlewood-Offord problem can be solved for Bernoulli distribution of arbitrary parameter $p\in(0,1)$. This question has been investigated by Singhal [24], where the author provides partial answers. A full solution was given by Melbourne, Madiman and Roberto in [16].
@@
The goal of this article is to extend the solution of the Littlewood-Offord problem to the whole class of discrete log-concave distributions. In particular, we prove the following.
@@
$\sup_{x\in\mathbb{R}}\,\mathbb{P}(a_{1}X_{1}+\cdots+a_{n}X_{n}\in A_{l,m}(x))\leq\frac{l}{\sqrt{1+c\sum_{k=1}^{n}{\rm Var}(X_{k})+c\,\frac{l^{2}-1}{12}}},$
@@
$\sup_{x\in\mathbb{R}}\,\mathbb{P}(a_{1}X_{1}+\cdots+a_{n}X_{n}\in A_{l,m}(x))\leq\frac{l}{\sqrt{1+\frac{n}{2}+\frac{l^{2}-1}{6}}}.$
We refer to Section 4 for an estimate when the $X_{k}$’s have a Bernoulli distribution with arbitrary parameter $p\in(0,1)$.
@@
The article is organized as follows. Section 2 provides the necessary background on the notion of majorization and rearrangement inequalities, which are the main ingredients in the proofs of Theorems 1.1 and 1.2. The proofs of Theorems 1.1 and 1.2 are postponed to Section 3. Section 4 presents a variant of the Littlewood-Offord problem for arithmetic progressions, in particular, Theorem 1.3 is proved. The last section focuses on the Bernoulli and uniform distributions, where Theorem 1.4 is proved.
@@ 3 The Littlewood-Offord problem for discrete log-concave distributions and an entropic version
The first step in establishing the Littlewood-Offord problem for log-concave distributions is to reduce the problem to signs.
@@
Let us first observe that $a\cdot X$ can be majorized by $\tilde{a}\cdot X$, for set of constants $\tilde{a}_{i}\in\mathbb{Z}\setminus\{0\}$. Indeed, it has been observed in [16, Proof of Lemma 5.1] that one may construct a linear map $T\colon\mathbb{R}\to\mathbb{Q}$ such that $T(a_{i})\in\mathbb{Z}\setminus\{0\}$ for all $i$. Thus, by Lemma 2.2, $a\cdot X\prec T(a\cdot X)$. Further, since the $X_{i}$’s are integer-valued, one has
@@
Since $\alpha$-Rényi entropy is Schur concave as a consequence of [18, Proposition 3-C.1] (see also [17]), the proof of Theorem 1.2 follows immediately.
@@ 5.1 Bernoulli distribution
This section focuses on strengthening Theorem 1.3 for the Bernoulli distribution. Let $U_{l}$ be uniform on $\{1,\dots,l\}$. It has been shown in [20, Proof of Theorem 1.7] that for all $p\geq 2$ the Fourier transform of $U_{l}$ satisfies
@@
where $\Phi(x)=\frac{1}{\sqrt{x}}\int_{0}^{\sqrt{x}}e^{-t^{2}/2}dt$ and $c=\pi(l^{2}-1)A^{2}$. Note that
@@
where $\operatorname{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$ is the error function. On the other hand, it has been shown in [16, Theorem 2.8] that if $X$ is a Bernoulli random variable with variance $\sigma^{2}$ and $q\geq 1$,
@@
Using the bound $\Phi(z)\leq\frac{1}{\sqrt{1+\frac{z}{3}}}$ holding for all $z>0$ (see [16, Lemma 2.9]) we deduce
@@ 5.2 Discrete uniform distribution
where the first identity follows from Parseval. Therefore, using [16, Lemma 2.7], we obtain for $n$ independent uniformly distributed random variables $U_{1},\dots,U_{n}$

### v2

@@ 1 Introduction
Kleitman extended this result when $a_{1},\dots,a_{n}$ are vectors in a Hilbert space [16], [17]. Many variants of the Erdös-Littlewood-Offord problem have been established, such as improved bounds under certain constraints on the $a_{i}$’s [11], an inverse Littlewood-Offord theorem [30], [26], and a resilience version [2], but most are mainly dealing with Bernoulli distribution with parameter $1/2$. Recently, Fox, Kwan and Sauermann [9, Question 6.2] asked whether the Littlewood-Offord problem can be solved for Bernoulli distribution of arbitrary parameter $p\in(0,1)$. This question has been investigated by Singhal [28], who gave a qualitative solution to the problem, showing that a maximizer is obtained for $a_{i}\in\{-1,1\}$ and gave Fourier theoretic bounds. Sharp quantitative bounds were found by Madiman, Melbourne, and Roberto in [18] where an entropic generalization of the problem was considered. The case of general random variables was considered recently by Juškevičius and Kurauskas [15].
@@
The goal of this article is to extend the solution of the Littlewood-Offord problem to the whole class of discrete log-concave distributions. Our bounds are quantitative and non-asymptotic. In particular, we prove the following.
@@
$\sup_{a\in(\mathbb{R}\setminus\{0\})^{n}}\sup_{x\in\mathbb{R}}\,\mathbb{P}(a_{1}X_{1}+\cdots+a_{n}X_{n}\in A_{l,m}(x))\leq\frac{l}{\sqrt{1+c\sum_{k=1}^{n}{\rm Var}(X_{k})+c\,\frac{l^{2}-1}{12}}},$
@@
$\sup_{a\in(\mathbb{R}\setminus\{0\})^{n}}\sup_{x\in\mathbb{R}}\,\mathbb{P}(a_{1}X_{1}+\cdots+a_{n}X_{n}\in A_{l,m}(x))\leq\frac{l}{\sqrt{1+\frac{n}{2}+\frac{l^{2}-1}{6}}}.$
We refer to Section 5 for an estimate that improves Theorem 1.3 when the $X_{k}$’s have a Bernoulli distribution with arbitrary parameter $p\in(0,1)$, $p\neq\frac{1}{2}$.
@@
The article is organized as follows. Section 2 provides the necessary background on the notion of majorization and rearrangement inequalities, which are the main ingredients in the proofs of Theorems 1.1 and 1.2. The proofs of Theorems 1.1 and 1.2 are postponed to Section 3. Section 4 demonstrates the advantage of our results compared with general bounds on the concentration function existing in the literature. Section 5 presents a variant of the Littlewood-Offord problem for arithmetic progressions, in particular, Theorem 1.3 is proved. The last section focuses on the Bernoulli and uniform distributions, where Theorem 1.4 is proved.
@@ 3 The Littlewood-Offord problem for discrete log-concave distributions and an entropic version
Throughout this section, given random variables $X_{1},\dots,X_{n}$, we denote $X=(X_{1},\dots,X_{n})$. The first step in establishing the Littlewood-Offord problem for log-concave distributions is to reduce the problem to signs.
@@
Let us first observe that $a\cdot X$ can be majorized by $\tilde{a}\cdot X$, for set of constants $\tilde{a}_{i}\in\mathbb{Z}\setminus\{0\}$. Indeed, it has been observed in [18, Proof of Lemma 5.1] that one may construct a linear map $T\colon\mathbb{R}\to\mathbb{Q}$ such that $T(a_{i})\in\mathbb{Z}\setminus\{0\}$ for all $i$. Thus, by Lemma 2.2, $a\cdot X\prec T(a\cdot X)$. Further, since the $X_{i}$’s are integer-valued, one has
@@
Since $\alpha$-Rényi entropy is Schur concave as a consequence of [20, Proposition 3-C.1] (see also [19]), the proof of Theorem 1.2 follows immediately.
@@
4 Comparison with general bounds on the concentration function
The goal of this section is to demonstrate that our upper bound on
$M(X)=\sup_{x\in\mathbb{R}}\mathbb{P}(X=x)$
given in equation (27), specialized to $a=(1,\dots,1)$, provides better information compared with existing results in the literature on concentration function. Let us recall that the concentration function of a random variable $X$ is defined as
$Q(X;\lambda)=\sup_{x}\mathbb{P}(x\leq X\leq x+\lambda),\qquad\lambda\geq 0.$
Note that for $X$ integer valued and $\lambda<1$,
$Q(X;\lambda)=M(X).$
A general bound was established by Miroshnikov and Rogozin [25].
Theorem 4.1 (Miroshnikov-Rogozin [25]).
There exists a universal constant $C>0$ such that $2\lambda\geq\lambda_{k}$ and $X_{k}$ independent random variables with $S=X_{1}+\cdots+X_{n}$,
$Q(S;\lambda)\leq C\lambda\left(\sum_{k=1}^{n}\mathbb{E}\left(|X_{k}^{s}|\wedge\frac{\lambda_{k}}{2}\right)^{2}Q^{-2}(X_{k},\lambda_{k})\right)^{-\frac{1}{2}},$
where given $X$, $X^{s}=X-X^{\prime}$ and $X^{\prime}$ is an independent copy of $X$.
[…]
$M(S)=Q(S;\lambda)\leq C\lambda\left(\sum_{k=1}^{n}\frac{\lambda_{k}^{2}\,2p_{k}(1-p_{k})}{4}M^{-2}(X_{k})\right)^{-\frac{1}{2}}.$
Minimizing the right hand side with $\lambda_{k}=2\lambda$ we have
$M(S)\leq C\left(\sum_{k=1}^{n}\frac{2p_{k}(1-p_{k})}{(p_{k}\vee(1-p_{k}))^{2}}\right)^{-\frac{1}{2}}=C\left(\sum_{k=1}^{n}\frac{2{\rm Var}(X_{k})}{(p_{k}\vee(1-p_{k}))^{2}}\right)^{-\frac{1}{2}}.$ (29)
Since $\frac{1}{2}\leq p\vee(1-p)\leq 1$ for $p\in[0,1]$, the right-hand side of (29) is at least $\frac{C}{\sqrt{2}}\left(\sum_{k=1}^{n}{\rm Var}(X_{k})\right)^{-\frac{1}{2}}$. Therefore for Bernoulli distributions, even after optimizing, the Miroshnikov-Rogozin yields the bound
$M(S)\leq\frac{\widetilde{C}}{\sqrt{{\rm Var}(S)}}$
for some absolute constant $\widetilde{C}>0$.
Note that in the absence of an explicit constant $\widetilde{C}$, this inequality is only interesting in the regime that ${\rm Var}(S)\to\infty$, where central limit theorems are often more viable. In particular, it yields trivial results in the Poisson regime, namely $M(S)\leq O(1)$, when the variance of $S$ can be bounded away from zero. In regimes where the variance tends to 0, the result is of an order worse than the trivial $M(S)\leq 1$. Whereas, our bounds derived in Section 3 provide meaningful quantitative estimates in all regimes.
In more recent work, explicit constants have been obtained for a variant of the Miroshnikov-Rogozin inequality. To this end, for $\theta$ a unit vector in $\mathbb{R}^{n}$ define
$p_{\theta}(t)=\mathrm{Vol}_{n-1}\left\{x\in\mathbb{R}^{n}:\|x\|_{\infty}\leq\frac{1}{2},\langle x,\theta\rangle=t\right\}.$
Theorem 4.2 (Bobkov-Chistyakov [3]).
Given $\lambda\geq\left(\sum_{k=1}^{n}\lambda_{k}^{2}\right)^{\frac{1}{2}}$ and $X_{k}$ independent random variables with $S=X_{1}+\cdots+X_{n}$,
$Q(S;\lambda)\leq\frac{2^{\frac{3}{2}}\lambda}{c}\left(\sum_{k=1}^{n}\lambda_{k}^{2}Q^{-2}(X_{k};\lambda_{k})\right)^{-\frac{1}{2}},$
where
$c\coloneqq\inf_{|t|<\frac{1}{2},\|\theta\|=1}p_{\theta}(t).$
Moreover, Bobkov and Chistyakov showed that $c\geq 0.00095$ independent of dimension. Later Melbourne, Tkocz, and Wyczesany [24] showed that the body $\{x\in\mathbb{R}^{n}:\|x\|_{\infty}\leq\frac{1}{2}\}$ can be replaced by any isotropic convex body $K$. More precisely for a convex body $K$ such that $\int_{K}x_{i}x_{j}dx=L_{K}^{2}\delta_{ij}$, for some constant $L_{K}>0$,
$\inf_{\theta,|t|<L_{K}\sqrt{3}}\mathrm{Vol}_{n-1}\left\{x\in K:\langle x,\theta\rangle=t\right\}\geq\frac{e^{-\sqrt{6}}}{\sqrt{2L_{K}^{2}}}.$
This result is proven sharp for high dimensional double cones, but even in the case of the cube where $L_{K}=\frac{1}{\sqrt{12}}$ it gives $c\geq\sqrt{6}e^{-\sqrt{6}}$, so that Bobkov and Chistyakov can be stated as
$Q(S;\lambda)\leq\frac{2e^{\sqrt{6}}}{\sqrt{3}}\lambda\left(\sum_{k=1}^{n}\lambda_{k}^{2}Q^{-2}(X_{k};\lambda_{k})\right)^{-\frac{1}{2}}.$
For reference, numerically $\frac{2e^{\sqrt{6}}}{\sqrt{3}}\ \approx 13.3742$. Hence for $C=\frac{2e^{\sqrt{6}}}{\sqrt{3}}$, $X_{k}$ integer valued and $\sum_{k=1}^{n}\lambda_{k}^{2}\leq\lambda^{2}<1$, by Bobkov-Chistyakov
$M(S)=Q(S;\lambda)\leq C\lambda\left(\sum_{k=1}^{n}\lambda_{k}^{2}Q^{-2}(X_{k};\lambda_{k})\right)^{-\frac{1}{2}}=C\lambda\left(\sum_{k=1}^{n}\lambda_{k}^{2}M^{-2}(X_{k})\right)^{-\frac{1}{2}}.$
Minimizing the right-hand side over $\lambda_{k}$, that is when $\lambda_{j}=\lambda$ for $j$ such that $M^{-2}(X_{j})=\max_{1\leq k\leq n}M^{-2}(X_{k})$ and $\lambda_{l}=0$ for $l\neq j$, gives
$M(S)\leq C\left(\max_{1\leq k\leq n}M^{-2}(X_{k})\right)^{-\frac{1}{2}}=C\min_{1\leq k\leq n}M(X_{k}).$
However, this is a trivial bound as Young’s convolution inequality yields
$M(S)\leq\min_{1\leq k\leq n}M(X_{k}).$
@@ 5 A Littlewood-Offord type problem for arithmetic progressions
Remark 5.1.
Let us note that Theorem 1.3 is sharp up to an absolute constant when taking supremum over all $m\neq 0$, as
$\sup_{m\neq 0}\sup_{a\in(\mathbb{R}\setminus\{0\})^{n}}\sup_{x\in\mathbb{R}}\,\mathbb{P}(a_{1}X_{1}+\cdots+a_{n}X_{n}\in A_{l,m}(x))\geq\frac{l}{\sqrt{1+12\sum_{k=1}^{n}{\rm Var}(X_{k})+l^{2}-1}},$
due to the lower bound in (5).
@@ 6.1 Bernoulli distribution
This section focuses on strengthening Theorem 1.3 for the Bernoulli distribution. Let $U_{l}$ be uniform on $\{1,\dots,l\}$. It has been shown in [22, Proof of Theorem 1.7] that for all $p\geq 2$ the Fourier transform of $U_{l}$ satisfies
@@
where $\Phi(x)\coloneqq\frac{1}{\sqrt{x}}\int_{0}^{\sqrt{x}}e^{-t^{2}/2}dt$ and $c=\pi(l^{2}-1)A^{2}$. Note that
@@
where $\operatorname{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$ is the error function. On the other hand, it has been shown in [18, Theorem 2.8] that if $X$ is a Bernoulli random variable with variance $\sigma^{2}$ and $q\geq 1$,
@@
Using the bound $\Phi(z)\leq\frac{1}{\sqrt{1+\frac{z}{3}}}$ holding for all $z>0$ (see [18, Lemma 2.9]) we deduce
@@ 6.2 Discrete uniform distribution
where the first identity follows from Parseval. Therefore, using [18, Lemma 2.7], we obtain for $n$ independent uniformly distributed random variables $U_{1},\dots,U_{n}$